### v1

# Computable Følner sequences of amenable groups

Abstract
The paper considers computable Følner sequences in computably enumerable amenable groups, We extend some basic results of M. Cavaleri on existence of such sequences to the case of groups where finite generation is not assumed. We also initiate some new directions in this topic, for example complexity of families of effective Følner sequences. Possible extensions of this approach to metric groups are also discussed.
1 Introduction
Analysis of classical mathematical topics from the point of view of complexity of various types is one of the major trends of modern mathematical logic. Amenability is essentially fruitful from this point of view (see [1], [6],[18], [16], [17], [20], [21], [25] ). Our research belongs to computable amenability. This is a topic where computable versions of fundamentals of amenability are studied, see papers of M. Cavaleri [3, 4], N. Moryakov [27] and the authors [8, 9]. Let us also mention [33] initiating a very rich field where computability meets topological dynamics.
In the present paper we return to the results of M. Cavaleri from [4], which are now considered as the beginning of the topic. It has been shown in [4] that amenable finitely generated recursively presented groups have computable Reiter functions and subrecursive Følner functions. Furthermore, for such a group decidability of the word problem is equivalent to so called effective amenability, i.e. existence of an algorithm which finds $\frac{1}{n}$-Følner sets for all $n$.
Since being finitely generated is not necessary for amenability, the question arises what happens if we consider the case of recursively presented groups without the assumption of finite generation. According the approach of computable algebra, the question concerns the class of computably enumerable numbered groups and the subclass of computable numbered groups, a counterpart of decidability of the word problem. These notions are thoroughly discribed in Section 2. The following theorem generalizes aforementioned results of Cavaleri to the case of computably enumerable numbered groups.
Theorem 1.
Let $(G,\nu)$ be a computably enumerable group. The following conditions are equivalent:
  1. (i)
    $(G,\nu)$ is amenable;
  2. (ii)
    $(G,\nu)$ has computable Reiter functions;
  3. (iii)
    $(G,\nu)$ has subrecursive Følner function.
  4. (iv)
    $(G,\nu)$ is $\Sigma$-amenable (see Definition 3.1).
Furthermore, computable amenability of $(G,\nu)$ is equivalent to computability of it.
This theorem summarizes our results of Section 3.
In the second part of the paper (Section 4) we concentrate on algorithmic complexity of effective Følner sequences and families of these sequences. In particular, we prove the following theorem.
Theorem 2.
The the set of all effective Følner sequences of a computable group belongs to the class $\Pi^{0}_{3}$, and, furthermore, in some cases of abelian groups this family is $\Pi^{0}_{3}$-complete.
We also compare convergence moduli of sequences of means corresponding to these Følner sequences. In particular we show that they cannot be estimated by primitive recursive functions.
In the final part of our paper (Section 5) we study possible generalizations of our results to computable metric groups. We suggest a framework to computable amenability in this general case. In particular, we define and discuss counterparts of basic notions studied in the main body of the paper. In these terms we prove the following theorem.
Theorem 3.
A computably enumerable numbered metric group $(G,d,\nu)$ is computably amenable if and only if it is amenable and computable.
This is an extension of the final statement of Theorem 1 to the case of metric groups.
The author is grateful to M. Cavaleri and T. Ceccherini-Silberstein for reading of some preliminary versions of the paper and helpful remarks.
2 Preliminaries
In this section we give preliminaries on (a) computable groups and (b) amenable groups. While (b) is rather standard, the topic (a) is presented in the form which seems new. We have found that the notion of enumerated groups from [14] allows us to simplify the presentation of the material and some arguments.
Concerning other details, let us mention that we often identify finite sets $F\subset\mathbb{N}$ with their Gödel numbers. For any sets $X$ and $Y$ we will write $X\subset_{fin}Y$ to denote that $X$ is a finite subset of $Y$. For any $i\in\mathbb{N}$, we denote the set $\{1,2,\ldots,i\}$ by $[i]$. Throughout this paper a group (metric group) $G$ is a countable/separable group without any presumption about its generating set.
A function is subrecursive if it admits a computable total upper bound. A sequence $(n_{i})_{i\in\mathbb{N}}$ of natural numbers is called computable/effective, if the function $k\rightarrow n_{k}$ is recursive. We use standard material from the computability theory (see [34]) and often say computable instead of recursive.
2.1 Computabie presentations
Let $G$ be a countable group generated by some $X\subseteq G$. The group $G$ is called recursively presented (see Section IV.3 in [24]) if $X$ can be identified with $\mathbb{N}$ (or with some $\{0,\ldots,n\}$) so that $G$ has a recursively enumerable set of relators in $X$. Below we give an equivalent definition, see Definition 2.2. It is justified by a possibility identification of the whole $G$ with $\mathbb{N}$. We develope the approaches of [11], [12] and [22].
Definition 2.1.
Let $G$ be a group and $\nu:\mathbb{N}\rightarrow G$ be a surjective function. We call the pair $(G,\nu)$ a numbered group. The function $\nu$ is called a numbering of $G$. If $g\in G$ and $\nu(n)=g$, then $n$ is called a number of $g$.
In this paper we usually assume that numberings of the group are homomorphisms from some group defined on $\mathbb{N}$. This condition is formulated in the following definition. The notion of enumerated group used in it, is taken from [14].
Definition 2.2.
  • A group of the form $(\mathbb{N},\star,^{-1},1)$ (where the number $1$ is the neutral element of the group) is called an enumerated group.
  • Given an enumerated group $(\mathbb{N},\star,^{-1},1)$ we call a surjective homomorphism $\nu:\mathbb{N}\to G$ a computably enumerable presentation of $G$ if the set
[…]
    $w(\nu(n_{1}),\ldots,\nu(n_{s}))=w^{\prime}(\nu(\ell_{1}),\ldots,\nu(\ell_{t}))\mbox{ holds in }G\,\,,\,n_{1},\ldots,n_{s},\ell_{1},\ldots,\ell_{t}\in\mathbb{N}\}$
    is computably enumerable.
An easy folklore argument shows that every finitely generated group with decidable word problem can be presented as an enumerated group $(\mathbb{N},\star,^{-1},1)$ such that $\star$ and the corresponding -1 are computable functions. This also holds in the case of the free group $\mathbb{F}_{\omega}$ with the free basis $\omega=\{0,\ldots,i,\ldots\}$. From now on let us fix such a presentation of $\mathbb{F}_{\omega}$:
$(\mathbb{N},*,^{-1},1).$
We emphasize that this group is computable! For every recursively presented group $G=\langle X\rangle$ and a natural homomorphism $\rho:\mathbb{F}_{\omega}\rightarrow G$ (taking $\omega$ onto $X$) we obtain that $\rho$ is a numbering of a computably enumerable presentation of $G$. This argument can be generalized as follows.
Lemma 2.3.
Assume that $(G,\nu)$ is a numbered group and the set $\mathsf{Wrd}^{=}_{\nu}$ is computably enumerable.
Then $G$ has a computably enumerable presentation $\nu^{0}:(\mathbb{N},*,^{-1},1)\to G$.
Proof. Note that extending the map $\nu$ to the set of all group words over the base $\omega=\{0,1,\ldots,i,\ldots\}$ we obtain a homomorphism $\rho:\mathsf{F}_{\omega}\to G$ such that the set
$\mathsf{Wrd}^{=}_{\rho}:=\{(w(u_{1},\ldots,u_{s}),w^{\prime}(v_{1},\ldots,v_{t}))\,|\,w(\bar{x})\mbox{ and }w^{\prime}(\bar{y})\mbox{ are group words, }$
$\bar{u},\bar{v}\in\mathbb{F}_{\omega}\mbox{ and the equation }w(\rho(u_{1}),\ldots,\rho(u_{s}))=w^{\prime}(\rho(v_{1}),\ldots,\rho(v_{t}))\mbox{ holds in }G\,\}$
is computably enumerable.
Under the coding, which we mentioned above, of the group $\mathbb{F}_{\omega}$ into the enumerated group $(\mathbb{N},*,^{-1},1)$ the homomorphism $\rho$ becomes the numbering $\nu^{0}:(\mathbb{N},*,^{-1},1)\to G$ as in the formulation. $\Box$
In particular, every group with a computably enumerable presentation has a presentation as in the lemma, i.e. where the enumerated group is computable. It is convenient to use the following notions too.
Definition 2.4.
  • If $G$ has a computably enumerable presentation then we say that $G$ is computably enumerable.
  • If a homomorphism $\nu:\mathbb{N}\to G$ is a computably enumerable presentation of $G$ and the set $\mathsf{Wrd}^{=}_{\nu}$ is computable (i.e. decidable), then we say that $\nu$ is a computable presentation and the group $G$ is computable.
Note that under the conditions of Lemma 2.3 the sets $\{n:\nu^{0}(n)=1\}$ and $\{(n_{1},n_{2}):\nu^{0}(n_{1})=\nu^{0}(n_{2})\}$ are computably enumerable. In the following remark we consider the case when they are computable.
Remark 2.5.
Let $\nu:(\mathbb{N},\star,^{-1},1)\to G$ be a computably enumerable presentation such that $\star$ and the corresponding -1 are computable functions.
  1. (i)
    The set
    $\mathsf{MultT}:=\{(i,j,k):\quad\nu(i)\nu(j)=\nu(k)\}$
    is computable if and only if the set $\{(n_{1},n_{2}):\nu(n_{1})=\nu(n_{2})\}$ is computable;
  2. (ii)
    If $\mathsf{MultT}$ is computable, the presentation $\nu$ is computable too and, furthermore, it can be made a computable bijective presentation.
Indeed, in this case the set of the smallest numbers of the elements of $G$ is computable. Enumerating this set by natural numbers we obtain a required bijective numbering.
Computable groups correspond to groups with solvable word problem. In this case the numbering $\nu$ is often called a constructivization, see [11], [12].
From now on we consider computably enumerable groups together with presentations satisfying the conditions of Lemma 2.3.
2.2 Amenability
The preliminaries of amenability correspond to [5] and [29]. Let $G$ be a group, and $\ell^{\infty}(G)$ be the Banach space of bounded functions $G\to\mathbb{R}$ with respect to the the norm
$\parallel{\bf x}\parallel_{\infty}=\mathsf{sup}\{|{\bf x}(a)|\,|\,a\in G\}.$
The group $G$ is called amenable if there is a left-invariant (equivalently rigth-invariant, resp. bi-invariant) mean $\ell^{\infty}(G)\to\mathbb{R}$ on $G$.
Given $n\in\mathbb{N}$ and $D\subset_{fin}G$, a subset $F\subset_{fin}G$ is called an $\frac{1}{n}$-Følner set with respect to $D$ if
$\forall x\in D\hskip 14.22636pt\frac{|F\setminus xF|}{|F|}\leq\frac{1}{n}$ (1)
We denote by $\mathfrak{F}$$l_{G,D}(n)$ the set of all $\frac{1}{n}$-Følner sets with respect to $D$. Moreover, we call the binary function:
$F\o l_{G}(n,D)=\mathsf{min}\{|F|\,|\,F\subseteq G\text{ such that }F\in\mathfrak{F}\o l_{G,D}(n)\},$ (2)
where the variable $D$ corresponds to finite sets, the Følner function of $G$.
A sequence $(F_{j})_{j\in\mathbb{N}}$ of non-empty finite subsets of $G$ is a Følner sequence if for every $g\in G$ the following condition holds:
$\lim\limits_{j\rightarrow\infty}\frac{|F_{j}\setminus gF_{j}|}{|F_{j}|}=0.$ (3)
It is easy to see that existence of Følner sets for all $n$ and $D$ is equivalent to existence of a Følner sequence:
$\bullet$ $G$ admits a Følner sequence if and only if $F\o l_{G}(n,D)<\infty$ for all $n\in\mathbb{N}$ and $D\subset_{fin}G$.
In fact, this is the Følner condition of amenability.
The following example will be helpful below. The group $(\mathbb{Z},+)$ has the following Følner sequence:
$\mathcal{F}=(\{-i,-i+1,\ldots,0,\ldots,i-1,i\}\,|\,i\in\mathbb{N}).$
Note that $\{-i,-i+1,\ldots,0,\ldots,i-1,i\}$ is $\frac{1}{2i}$-Følner with respect to the generator of $\mathbb{Z}$.
Suppose that $G$ satisfies the Følner conditions. Let $(F_{j}\,|\,j\in J)$ be a left Følner’s net of $G$. Consider, for each $j\in J$, the mean with finite support $m_{j}:\ell^{\infty}(G)\to\mathbb{R}$ defined by
$m_{j}({\bf x})=\frac{1}{|F_{j}|}\sum\{{\bf x}(h)\,|\,h\in F_{j}\}\mbox{ for all }{\bf x}\in\ell^{\infty}(G).$
The net $(m_{j}\,|\,j\in J)$ contains a sequence converging (in the weak topology) to an invariant mean $m$ (see Lemma 4.5.9 and Theorem 4.9.2 in [5]).
The space of absolutely summable functions $\ell^{1}(G)$ is considered with respect to the norm
$\parallel{\bf x}\parallel_{1}=\sum\{|{\bf x}(a)|\,|\,a\in G\}\mbox{ , where }{\bf x}:G\to\mathbb{R}\mbox{ is absolutely summable.}$
Definition 2.6.
A non-zero function $h:G\rightarrow\mathbb{R}^{+}$, $h||_{1}<\infty$, is $\frac{1}{n}$-invariant with respect to $D$, if
[…]
where $_{x}h(g):=h(x^{-1}g)$.
We denote by $\mathfrak{R}eit_{G,D}(n)$ the set of all summable non-zero functions from $G$ to $\mathbb{R}^{+}$, which are $\frac{1}{n}$-invariant with respect to $D$.
The following facts are well known and/or easy to prove.
Lemma 2.7.
Let $F,D\subset_{fin}G$.
  1. (i)
    $F\in\mathfrak{F}\o l_{G,D}(n)\implies\forall g\in G\quad Fg\in F\o l_{G,D}(n)$
  2. (ii)
    $F\in\mathfrak{F}\o l_{G,D}(n)\iff\forall x\in D\quad\frac{|F\cap xF|}{|F|}>1-\frac{1}{n}$
  3. (iii)
    $F\in\mathfrak{F}\o l_{G,D}(2n)\iff\chi_{F}\in\mathfrak{R}eit_{G,D}(n)$
  4. (iv)
    If $h\in\mathfrak{R}eit_{G,D}(n)$ has finite support then there exists $F\subset\mathsf{supp}(h)$ such that for all $x\in D$ the following holds:
    $\frac{|F\setminus xF|}{|F|}<\frac{|D|}{2n}.$
Let $Prob(G)$ be the set of countably supported probability measures on $G$ (with respect to the $\sigma$-algebra of all subsets of $G$). Under the assumption (in Sections 3 and 4) that $G$ is countable, $Prob(G)$ is exactly the set of probability measures on $G$. Any element of this space can be written $\mu=\sum_{i\in\mathbb{N}}\lambda_{i}\chi_{g_{i}}$ where $g_{i}\in G$, $\lambda_{i}\geq 0$ and $\sum_{i\in\mathbb{N}}\lambda_{i}=1$. In particular, it can identified with a subset of the unit ball of $\ell^{1}(G)$.
Reiter’s condition states that the group is amenable if and only if for any finite subset $D\subset G$ and $\varepsilon>0$, there is $\mu\in Prob(G)$ such that $\parallel\mu-_{g}\mu\parallel_{1}\leq\varepsilon$ for any $g\in D$. This justifies the term $\mathfrak{R}eit_{G,D}(n)$.
3 Effective amenability of computably enumerable groups
In this section $(G,\nu)$ is a numbered metric group such that $\nu$ is a homomorphism from $(\mathbb{N},*,^{-1},1)$ to $G$. In the case of the Følner condition of amenability, we consider two types of effectiveness.
Definition 3.1.
The numbered group $(G,\nu)$ is $\Sigma$-amenable, if there is an algorithm which for all pairs $(n,D)$ where $n\in\mathbb{N}$ and $D\subset_{fin}\mathbb{N}$, finds a set $F\subset_{fin}\mathbb{N}$ having a subset $F^{\prime}\subseteq F$ with $\nu(F^{\prime})\in\mathfrak{F}\o l_{G,\nu(D)}(n)$.
Definition 3.2.
The numbered group $(G,\nu)$ is computably amenable if there exists an algorithm which for all pairs $(n,D)$, where $n\in\mathbb{N}$ and $D\subset_{fin}\mathbb{N}$, finds a finite set $F\subset\mathbb{N}$ such that $\nu(F)\in\mathfrak{F}\o l_{G,\nu(D)}(n)$ and $|F|=|\nu(F)|$.
These are the main notions of our paper. It is clear that computable amenability implies $\Sigma$-amenability. Furthermore, $\Sigma$-amenability implies that the Følner function is subrecursive.
Some variants of these notions were characterized by M. Cavaleri in [4] in the case of finitely generated groups. The goal of this section is an adaptation of these characterizations in our general case.
3.1 Computable Reiter’s functions
The main result of this part, Theorem 3.5, is a natural generalization of a theorem of M. Cavaleri from [4] (Theorem 3.1) to the case of groups which are not finitely generated. Although we use the same arguments, its adaption to our case is not straightforward. Throughout this section we assume that $(G,\nu)$ is a computably enumerable group under a homomrphism $\nu$ as in the beginning of the section.
The following notation will be used below. If $f:\mathbb{N}\to\mathbb{R}^{+}$ is summable and $g\in G$, then let
$\nu_{G,1}(f)(g)=\sum\{f(i)\,|\,i\in\nu^{-1}(g)\}.$
Definition 3.3.
We say that $(G,\nu)$ has computable Reiter functions, if there exists an algorithm which, for every $n\in\mathbb{N}$ and any finite set $D\subset\mathbb{N}$ finds $f:\mathbb{N}\rightarrow\mathbb{Q^{+}}$, such that $|\mathsf{supp}(f)|<\infty$ and
[…]
We now need some preliminary material concerning Reiter functions and partitions. Let $X$ be a nonempty set. An equivalence relation $E^{\prime}$ on $X$ is called finer than an equivalence relation $E$ if $E^{\prime}\subseteq E$.
Let $f:\mathbb{N}\rightarrow\mathbb{Q}_{+}$ be a function with a finite support $F$ and let $X$ be a finite set including $F$. For an equivalence relation $E$ of $X$ define $E_{F,1}(f)$ as follows:
$E_{F,1}(f)(x):=\sum\{f(i)\,|\,i\in F\mbox{ and }(i,x)\in E\}.$
Having $x\in\mathbb{N}$ and an equivalence relation $E$ on the set $F\cup x^{-1}*F$ fix a set $F_{0}\subset F$ of representatives of the classes of $E$ in $F$. Then we define the positive rational number:
$M_{E,F}^{x}(f):=\frac{\sum\{|(E_{F,1}(f)(v)-E_{F,1}(f)(x^{-1}*v))|\,|\,v\in F_{0}\}}{\sum_{v\in F}f(v)}.$
We denote by $P$ the canonical equivalence relation of the set $F\cup x^{-1}*F$, i.e. the partition into sets $\{\nu^{-1}(\nu(k))\,|\,k\in F\cup x^{-1}*F\}$. Then for every $x\in\mathbb{N}$ we have
[…]
For any two equivalence relations $E$ and $E^{\prime}$ of the set $F\cup x^{-1}*F$ with $E\subseteq E^{\prime}$ the triangle inequality implies $M_{E,F}^{x}(f)\geq M_{E^{\prime},F}^{x}(f)$. In particular, for any $x\in\mathbb{N}$ and $E\subseteq P$ on $F\cup x^{-1}*F$ we have:
$M_{E,F}^{x}(f)\geq M_{P,F}^{x}(f).$ (6)
Lemma 3.4.
Let $(G,\nu)$ be a computably enumerable group as above. There exists a computable enumeration of the set of all triples $(n,D,f)$, where $D\subset_{fin}\mathbb{N}$ and $f:\mathbb{N}\rightarrow\mathbb{Q}^{+}$ is a finitely supported function, such that $\nu_{G,1}(f)\in\mathfrak{R}eit_{G,\nu(D)}(n)$.
Proof.
[…]
The algorithm $\varkappa(n,D,f)$ starts as follows. For an input $f$ let $F=\mathsf{supp}(f)$. Put $P_{0}$ to be the (finest) partition of $\bigcup_{g\in D}g^{-1}*F$ into singletons. At the $m$-th step of the enumeration of $\{(n_{1},n_{2}):\nu(n_{1})=\nu(n_{2})\}$ we are trying to merge classes of the equivalence relation $P_{m-1}$ on $\bigcup_{g\in D}g^{-1}*F$ already constructed at step m1m-1. We do so when we meet $(n_{1},n_{2})\not\in P_{m-1}$ with $\nu(n_{1})=\nu(n_{2})$. In this case we just merge the classes of $n_{1}$ and $n_{2}$. In this way we obtain $P_{m}\subseteq P$. Then we verify if $M_{P_{m},F}^{x}(f)\leq\frac{1}{n}$ for all $x\in D$. We stop $\varkappa(n,D,f)$ when these inequalities hold. In this case by (5) and (6) the function $\nu_{G,1}(f)$ is $\frac{1}{n}$-invariant. If there exist $x$, such that $M_{P_{m},F}^{x}(f)>\frac{1}{n}$ and $P_{m}=P$, then the function $\nu_{G,1}(f)$ is not $\frac{1}{n}$-invariant. Note that there is no algorithm for recognizing the latter possibility.
The algorithm stated in the formulation of the lemma at $k$-th step makes the first move in $\varkappa(n_{k},D_{k},f_{k})$, the second move in $(n_{k-1},D_{k-1},f_{k-1}),\ldots$, and the $k$-th move in $\varkappa(n_{1},D_{1},f_{1})$. When one of these procedures gives $\nu_{G,1}(f)\in\mathfrak{R}eit_{G,\nu(D)}(n)$ we put the corresponding triple into our list. ∎
The following theorem is a part of Therem 1 from the introduction. The proof uses the procedure $\varkappa(n,D,f)$ from the proof of Lemma 3.4.
Theorem 3.5.
Let $(G,\nu)$ be a computably enumerable group. Then the following conditions are equivalent:
  1. (i)
    $G$ is amenable;
  2. (ii)
    $(G,\nu)$ has a subrecursive Følner function;
  3. (iii)
    $(G,\nu)$ is $\Sigma$-amenable;
  4. (iv)
    $(G,\nu)$ has computable Reiter functions.
Proof.
[…]
3.2 Effective amenability of computable groups
The main results of this section, correspond to Theorem 4.1 and Corollary 4.2 of M. Cavaleri from [4]. In the proof we will use functions * and -1 from Lemma 2.3. In particular, we are under the conditions of Remark 2.5.
Theorem 3.6.
Let $(G,\nu)$ be a computably enumerable group. The following conditions are equivalent:
  1. (i)
    $(G,\nu)$ is amenable and computable;
  2. (ii)
    $(G,\nu)$ is computably amenable (Definition 3.2).
Proof.
[…]
Fix $n_{1},n_{2}$. Let $E$ be the set $\{n_{1},n_{2}\}$. We use the algorithm for (ii) to find a set $F$ corresponding to $3$ and $E$, i.e. $\nu(F)\in F\o l_{G,\nu(E)}(3)$ and $|F|=|\nu(F)|$. Let $F=\{f_{1},f_{2},\ldots,f_{k}\}$.
For each $i\in\{1,2\}$ we define $\Sigma_{i}\subseteq\{(f,f^{\prime})\,|\,\nu(n_{i})\nu(f)=\nu(f^{\prime})\,,\,f,f^{\prime}\in F\}$ by the following procedure. Having $f,f^{\prime}\in F$ apply the algorithm of enumeration of the set $\mathsf{Wrd}^{=}_{\nu}$ for verification if $\nu(n_{i}*f)=\nu(f^{\prime})$, $i=1,2$. When we get a confirmation of this equality, we extend the corresponding $\Sigma_{i}$ by $(f,f^{\prime})$. We apply it simultaneously to each pair $(f,f^{\prime})$. Since
$\forall n\in E\,\,(|\nu(F)\cap\nu(n)\nu(F)|\geq\frac{2}{3}|\nu(F)|)$
and $|F|=|\nu(F)|$, there is a step of these computations when $\Sigma_{1}\cup\Sigma_{2}$ witnesses the inequality above. Having this we stop the procedure.
[…]
The proof of Theorem 3.6 gives the following interesting observation.
Corollary 3.7.
Let $(G,\nu)$ be a computably enumerable, amenable group. If for some $n\geq 3$ there exists an algorithm, which for every $D\subset_{fin}\mathbb{N}$ with $|D|=2$, finds a set $F\subset_{fin}\mathbb{N}$ such that $\nu(F)\in F\o l_{G,\nu(D)}(n)$ and $|F|=|\nu(F)|$, then $G$ is computable.
Using Theorem 3.6 we deduce a version of Theorem 3.5 for computable groups. This finishes the proof of Theorem 1.
Theorem 3.8.
Let $(G,\nu)$ be a computable group. Then the following conditions are equivalent:
  1. (i)
    $(G,\nu)$ is amenable;
  2. (ii)
    $(G,\nu)$ is computably amenable;
  3. (iii)
    there exists an algorithm which, for all pairs $(n,D)$, where $n\in\mathbb{N}$ and $D\subset_{fin}\mathbb{N}$, finds a finite set $F\subset\mathbb{N}$ such that $\nu(F)\in\mathfrak{F}$$l_{G,\nu(D)}(n)$ (a weaker version of (ii));
  4. (iv)
    $(G,\nu)$ has computable Reiter functions;
  5. (v)
    $(G,\nu)$ has subrecursive Følner function.
Proof.
[…]
4 Effective Følner sequence
Let $(G,\nu)$ be a computable group. By Remark 2.5 we may assume that the function $\nu$ is injective. Therefore, we identify the set $\nu(\mathbb{N})$ with $\mathbb{N}$ and subsets $F$ of $\mathbb{N}$ with $\nu(F)\subseteq G$.
A computable Følner sequence of the group $(G,\nu)$ is a computable sequence $(n_{j})_{j\in\mathbb{N}}$ where each $n_{j}$ is a Gödel number of some $F_{j}$, such that $(F_{j})_{j\in\mathbb{N}}$ is a Følner sequence. When $(n_{j})_{j\in\mathbb{N}}$ is computable we will also say that $(F_{j})_{j\in\mathbb{N}}$ is a computable (or effective) Følner sequence.
In this section we consider two questions where Følner sequences are involved. Firstly we analyse algorithmic complexity of the set of all computable Følner sequences. The second part of the section concerns the procedure how an invariant mean is built from a Følner sequence. Having ${\bf x}\in\ell^{\infty}(G)$ the corresponding value of the mean is the limit of an effective sequence of rational numbers. We will study the complexity of the corresponding convergence modulus. Questions of this kind are quite natural in computability theory, see [23] and the discussion in the end of the section.
4.1 Algorithmic complexity of computable Følner sequences
In the previous section we have shown that amenability of $(G,\nu)$ is equivalent to computable amenability. Note that this is also equivalent to existence of an effective Følner sequence. Indeed, given $\ell$ we put $E=\{0,\ldots,\ell\}$ and using the algorithm for computable amenability, compute the Gödel number $n_{\ell}$ of some $F_{\ell}\in\mathfrak{F}\o l_{G,E}(\ell)$. Clearly, the sequence $(F_{j})_{j\in\mathbb{N}}$ is a Følner’s one and the sequence $(n_{j})_{j\in\mathbb{N}}$ is a computable Følner sequence.
The following theorem classifies the set of all computable Følner sequences of the group $(G,\nu)$ in the arithmetical hierarchy. This is Theorem 2 from Introduction.
Theorem 4.1.
Let $(G,\nu)$ be a computable group. The set of all computable Følner sequences of $(G,\nu)$ belongs to the class $\Pi^{0}_{3}$. Moreover, for $G=\bigoplus\limits_{n\in\omega}\mathbb{Z}$ it is a $\Pi^{0}_{3}$-complete set.
Proof.
Let $\varphi(x,y)$ be a universal recursive function, and $\varphi_{x}(y)=\varphi(x,y)$ be the recursive function with the number $x$. We identify computable Følner sequences with numbers of recursive functions which produce these sequences. The set of these numbers will be denoted by $\mathfrak{F}_{seq}(G)$.
It is straightforward that $m\in\mathfrak{F}_{seq}(G)$ if and only if the following formula holds:
$(\phi(m,y)$ $\text{ is a total function})\wedge(\forall x,n\in\mathbb{N})(\exists l)(\forall k,s)\Big{(}k>l\wedge(\phi(m,k)=s)$
[…]
Given number $s$ the inequality $\frac{|F\setminus x*F|}{|F|}<\frac{1}{n}$ can be verified effectively. Since the set of numbers of all total functions belongs to the class $\Sigma_{2}^{0}$ (see [34]), it is easy to see that the set of all $m$ which satisfy (9) is a $\Pi_{3}^{0}$ set. This proves the first part of the theorem.
We remind the reader that $W_{t}=\mathsf{Dom}\varphi_{t}$ is the computably enumerable set with a number $t$. The set $\overline{Cof}=\{e:\forall n\;W_{\varphi_{e}(n)}\text{ is finite}\}$, is known to be a $\Pi^{0}_{3}$-complete set ([34], p. 87). To prove the second part of the theorem, assume that $G=\bigoplus\limits_{n\in\omega}\mathbb{Z}$. Let us show that the set $\overline{Cof}$ is reducible to $\mathfrak{F}_{seq}(G)$.
We present $\bigoplus\limits_{n\in\omega}\mathbb{Z}$ as $\bigoplus\limits_{n\in\omega}\langle g_{n}\rangle$. We shall construct a sequence $\{F_{s}^{e}\,|\,e,s\in\mathbb{N}\}$ such that $e\in\overline{Cof}$ iff $\{F_{s}^{e}\,|\,s\in\mathbb{N}\}$ is a Følner sequence. For each $e$ let us fix a computable enumeration of the set $\{(n,x):x\in W_{\varphi_{e}(n)}\}$. We can assume that this enumeration is without repetitions.
For a given $s$, we use the enumeration of the set $\{(n,x):x\in W_{\varphi_{e}(n)}\}$ to find the element $(n_{s},x_{s})$ with the number $s$. For each $i=1,\ldots,s$ such that $i\neq n_{s}$ let $F_{s,i}=\{0,g^{\pm 1}_{i},g_{i}^{\pm 2},\ldots,g_{i}^{\pm s}\}$. For $i=n_{s}$ we put $F_{s,i}=\{0,g^{\pm 1}_{i}\}$. Let $F_{s}^{e}=\bigoplus\limits_{1}^{s}F_{s,i}$. Then in the former case $F_{s}^{e}$ is an $\frac{1}{2s}$-Følner set with respect to $g^{\pm 1}_{i}$ and in the latter case $F_{s}^{e}$ is not a $\frac{1}{4}$-Følner set with respect to $g_{i}$. This ends the construction. Note that it depends on $e$ and corresponds to a Turing machime which can be explicitly found.
Case 1. $e\notin\overline{Cof}$. There exists $n^{\prime}$ such that $W_{\varphi_{e}(n^{\prime})}$ is an infinite set. Therefore, there exist an increasing sequence $\{s_{i}\}$ and the number $i^{\prime}$ such that for all $i>i^{\prime}$, $F_{s_{i}}^{e}$ is not an $\frac{1}{4}$-Følner set with respect to $g_{n^{\prime}}$. Clearly, the number of the algorithm producing the sequence $\{F_{s}^{e}\,|\,s\in\mathbb{N}\}$ does not belong to the set of numbers of Følner sequences.
Case 2. $e\in\overline{Cof}$. For all $n$, $W_{\varphi_{e(n)}}$ is a finite set. Therefore, for all $n$, there exists the number $s^{\prime}$ such that for all $s>s^{\prime}$, $F_{s}^{e}$ is an $\frac{1}{2s}$-Følner set with respect to $g^{\pm 1}_{n}$. Thus, by an easy argument we see that it is a Følner sequence.
Since for every $e$ the number of the algorithm producing $\{F^{e}_{s}\}$ can be effectively found it follows that the set $\overline{Cof}$ is reducible to $\mathfrak{F}_{seq}(G)$, which completes the proof. ∎
4.2 Complexity of convergence moduli
Suppose that $G$ satisfies the Følner conditions. Then $G$ admitsa a left Følner net $(F_{j}\,|\,j\in J)$. Consider, for each $j\in J$, the mean with a finite support $m_{j}:\ell^{\infty}(G)\to\mathbb{R}$, defined by
$m_{j}({\bf x})=\frac{1}{|F_{j}|}\sum\{{\bf x}(h)\,|\,h\in F_{j}\}\mbox{ for all }{\bf x}\in\ell^{\infty}(G).$
Then for every $g\in G$ and for every ${\bf x}\in\ell^{\infty}(G)$ we have $\lim_{j}(gm_{j}-m_{j})({\bf x})=0$ (see the proof of Theorem 4.9.2 in [5]). Taking a subnet if necessary we may assume that $(m_{j}\,|\,j\in J)$ converges (in the weak topology) to an invariant mean $m$ (see Lemma 4.5.9 and Theorem 4.9.2 in [5]).
Assume that $(G,\nu)$ is a computable group with an injective $\nu$. Let $(n_{j})_{j\in\mathbb{N}}$ define an effective Følner sequence of the group $(G,\nu)$, where each $n_{j}$ is a Gödel number of some $F_{j}$, such that $(F_{j})_{j\in\mathbb{N}}$ is a Følner sequence. The corresponding mean $m$ can be viewed as a measure $2^{G}\,\to\,[0,1]$. Note that if ${\bf x}\in 2^{G}$ is computable (with respect to $\nu$), then the sequence $(m_{j}({\bf x})\,|\,j\in J)$ is computable and converges to $m({\bf x})$. The question which we study in this section concerns moduli of this convergence.
Remark 4.2.
Having a computable Følner sequence $\mathcal{F}$ and a computable ${\bf x}\in 2^{G}$ let us define:
$Mod_{\mathcal{F}}({\bf x})=\{(k,j)\,|\,\forall j^{\prime},j^{"}(j^{"}>j<j^{\prime}\to|m_{j^{\prime}}({\bf x})-m_{j^{"}}({\bf x})|<\frac{1}{k}\}.$
It is easy to see that the complement of $Mod_{\mathcal{F}}({\bf x})$ in $\mathbb{N}\times\mathbb{N}$ is computably enumerable. Furthermore, when $Mod_{\mathcal{F}}({\bf x})$ is computable, the function
$mod_{\mathcal{F},{\bf x}}:k\to\mathsf{min}\{j\,|\,(k,j)\in Mod_{\mathcal{F}}({\bf x})\}$
is computable too.
The main question which we study below concerns the growth of $mod_{\mathcal{F},{\bf x}}$ introduced in this remark. In order to simplify the situation assume that $m({\bf x})$ is a given rational number. In fact, we will consider the ”pointed” set
$Mod^{p}_{\mathcal{F}}({\bf x})=\{(k,j)\,|\,\forall j^{\prime}(j<j^{\prime}\to|m_{j^{\prime}}({\bf x})-m({\bf x})|<\frac{1}{k}\}.$
As above, the complement of $Mod^{p}_{\mathcal{F}}({\bf x})$ in $\mathbb{N}\times\mathbb{N}$ is computably enumerable.
We will consider the group $\mathbb{Z}$ (under some standard 1-1 enumeration) and the effective Følner family
$\mathcal{F}=(\{-i,-i+1,\ldots,0,\ldots,i-1,i\}\,|\,i\in\mathbb{N}).$
(As we already noted in Section 2, $\{-i,-i+1,\ldots,0,\ldots,i-1,i\}$ is $\frac{1}{2i}$-Følner with respect to the generator of $\mathbb{Z}$.) Our main result below shows that the growth of $mod_{\mathcal{F},{\bf x}}$ is not bounded by a primitive recursive function. We believe that our main construction can be adapted to many other computable groups. The construction is presented in the proof of the following theorem.
Theorem 4.3.
Let $f$ be a total computable function $\mathbb{N}\to\mathbb{N}$. Then there is a computable ${\bf x}_{0}\in 2^{\mathbb{Z}}$ such that with respect to the computable Følner sequence $\mathcal{F}$ we have $m_{i}({\bf x}_{0})\to 0$, $i\in\mathbb{N}$, and for every $k\in\mathbb{N}$ there is $j>f(k)$ such that $|m_{j}({\bf x}_{0})|\geq\frac{1}{k}$.
Proof.
We define ${\bf x}_{0}\in 2^{\mathbb{Z}}$ such that ${\bf x}_{0}(0)=1$ and ${\bf x}_{0}(i)={\bf x}(-i)$. Further details are given in the inductive procedure below.
We start at $k=3$. At inductive step $k$ assume that at step k1>1k-1>1 we have already defined some number $i_{k-1}$ and values ${\bf x}_{0}(i)$ for all $i$ with $i_{k-1}\leq i\leq i_{k-1}$. We also assume that
$\frac{\sum\{{\bf x}_{0}(i)|-i_{k-1}\leq i\leq i_{k-1}\}}{2i_{k-1}+1}<\frac{1}{k-1}\leq\frac{\sum\{{\bf x}_{0}(i)|-i_{k-1}\leq i\leq i_{k-1}\}}{2i_{k-1}-1}.$
To have this condition at step k1=2k-1=2 we may assume that $i_{2}=5$ , ${\bf x}_{0}(1)={\bf x}_{0}(2)={\bf x}_{0}(3)=0$ and ${\bf x}_{0}(4)={\bf x}_{0}(5)=1$. Note that it now implies
$\frac{1}{k}<\frac{\sum\{{\bf x}_{0}(i)|-i_{k-1}\leq i\leq i_{k-1}\}}{2i_{k-1}+1}.$
Find $i^{\prime}_{k-1}\geq i_{k-1}$ such that
$\frac{1}{k}\leq\frac{2f(k)+\sum\{{\bf x}_{0}(i)|-i_{k-1}\leq i\leq i_{k-1}\}}{2f(k)+2i^{\prime}_{k-1}+1}<\frac{1}{k-1}.$
Let $t_{k}$ be the minimal natural number $i^{\prime}_{k}$ such that
$\frac{2f(k)+\sum\{{\bf x}_{0}(i)|-i_{k-1}\leq i\leq i_{k-1}\}}{2f(k)+2t_{k}+1}<\frac{1}{k}.$
Put $i_{k}=f(k)+t_{k}$ and define
[…]
Note that $i_{k}$ and the values ${\bf x}_{0}(i)$, $i\leq i_{k}$, are found in an effective way. In particular $I=\{i_{k}\,|\,k\in\mathbb{N}\}$ is a computable set and the sequence
$m_{j}({\bf x}_{0})=\frac{\sum\{{\bf x}_{0}(i)|-j\leq i\leq j\}}{2j+1}\,,\,j\in\mathbb{N}$
is computable. Furthermore, by the choice of $i_{k}$ we see that $i_{k}-1\geq f(k)$ and $m_{i_{k}-1}({\bf x}_{0})\geq\frac{1}{k}$.
To see that $m_{j}({\bf x}_{0})\to 0$ note that for every $k$ and every $j$ with $i_{k-1}\leq j\leq i_{k}$ the inequality $m_{j}({\bf x}_{0})\leq\frac{1}{k-1}$ holds. ∎
Corollary 4.4.
There is a computable ${\bf x}_{0}\in 2^{\mathbb{Z}}$ such that with respect to the computable Følner sequence $\mathcal{F}$ we have $m_{i}({\bf x}_{0})\to 0$, $i\in\mathbb{N}$, and for every primitive recursive function $f$ there is $k\in\mathbb{N}$ and $j>f(k)$ such that $|m_{j}({\bf x}_{0})|\geq\frac{1}{k}$.
Proof.
We enumerate all primitive recursive functions $g_{1},g_{2},\dots$ and define a function $\hat{g}$$\mathbb{N}\rightarrow\mathbb{N}$ by the rule
$\hat{g}(n)=\overset{n}{\underset{i=1}{\sum}}\overset{n}{\underset{j=1}{\sum}}\,\,g_{i}(j),\hskip 5.69054ptn\in\mathbb{N}.$
Clearly, $\hat{g}$ is computable and for each $k$ there is $n_{0}$ such that $g_{k}(n)\leqslant\hat{g}(n)$ for all $n>n_{0}$. Applying theorem above to $\hat{g}$ we obtain the statement of the corollary. ∎
Remark 4.5.
Let $\Gamma$ be a finitely generated group, $S\subseteq\Gamma$ a finite symmetric generating set. According to [10] a Følner sequence in the (colored) Cayley graph $Cay(\Gamma,S)$ is a sequence $F=\{F_{1},F_{2},\ldots)$ of finite spanned subgraphs such that for all natural $r\geq 0$ all but finitely many of the $F_{n}$ are $r$-approximations. The latter means that there exists a subset $W$ of vertices of $F_{n}$ of size $(1-1/r)|F_{n}|$ such that for any $w\in W$ the $r$-neighborhood of $w$ is rooted isomorphic to the $r$-neighborhood of a vertex of the Cayley-graph of $\Gamma$ (as edge labeled graphs). The group $\Gamma$ is amenable if $Cay(\Gamma,S)$ has a Følner sequence. Note, that the family $\mathcal{F}$ considered in Theorem 4.3 is a Følner sequence in the Cayley graph of $\mathbb{Z}$ with respect to the generators $\pm 1$.
Connections with computability over reals
The material of this section is connected with the topic of complexiity over reals initiated by Ko and Friedman in [23]. Consider a countable amenable group $G$ which is computable under some 1-1 numbering. In fact, we will assume that $G$ is defined on $\mathbb{N}$. Let $\mathcal{F}=(F_{0},F_{1},\ldots)$ be a computable Følner sequence in $G$. Consider $G$ as the following disjoint union:
$F_{0}\dot{\cup}(F_{1}\setminus F_{0})\dot{\cup}(F_{2}\setminus(F_{0}\cup F_{1}))\dot{\cup}\ldots\,=\{g_{1},g_{2},\ldots\}$
with $F_{i}\setminus(F_{0}\cup\ldots\cup F_{i-1})=\{g_{\ell_{i-1}+1},\ldots,g_{\ell_{i}}\}$. We identify ${\bf x}\in 2^{G}$ with
$\frac{{\bf x}(g_{1})}{2}+\frac{{\bf x}(g_{2})}{4}+\ldots+\frac{{\bf x}(g_{i})}{2^{i}}+\ldots\,.$
Now the following function maps $[0,1]$ into $[0,1]$:
$m_{i}({\bf x})=\frac{1}{|F_{i}|}\sum\{{\bf x}(h)\,|\,h\in F_{i}\}.$
Since this function has finitely many values, according to Lemma 2.1 from [23] and example (iii) after it, it is a partial recursive function from $[0,1]$ to $[0,1]$. At this stage we remind the reader that according to Definition 2.1 of [23] intuitively the computation of a partial recursive real function $f$ is as follows. For given dyadic number $\bf x$} and and a natural number $n$, the Turing machine tries to find a dyadic rational number $d_{n}$ of length $n$ such that $d_{n}$ is close to $f({\bf x})$ up to $\frac{1}{2^{n}}$. During the computation, $\bf x$} is used as an oracle. As a result $f({\bf x})=\lim_{n}d_{n}$.
Let $m({\bf x})=\lim_{i}m_{i}({\bf x})$. This function is a measure representing the invariant mean defined in the first paragraph of this section. Is it a partial recursive function? The answer is ”no” by the following reason. By Theorem 2.2 of [23] a partial recursive function is continuous on its domain. Note that
$\bullet$ for every interval $(q,q^{\prime})\subseteq[0,1]$, for every real $\varepsilon>0$ and for every set $X\subseteq G$ that is represented by some $r\in(q,q^{\prime})$ (as an element of $2^{G}$) there is $X^{\prime}\subseteq G$ and $r^{\prime}\in(q,q^{\prime})$ which represents $X^{\prime}$ such that $|r-r^{\prime}|<\varepsilon$ but $|m(r)-m(r^{\prime})|>\frac{1}{2}$.
Indeed, assuming that ${\bf x}\in 2^{G}$ corresponds to $X$ and $r$, choose a number $n$ such that
$q<\frac{{\bf x}(g_{1})}{2}+\frac{{\bf x}(g_{2})}{4}+\ldots+\frac{{\bf x}(g_{n})}{2^{n}}<q^{\prime}\,\mbox{ and }\Big{|}\frac{{\bf x}(g_{n+1})}{2^{n+1}}+\frac{{\bf x}(g_{n+2})}{2^{n+2}}+\ldots+\frac{{\bf x}(g_{i})}{2^{i}}+\ldots\,\Big{|}<\varepsilon.$
Then defining $\bf x$$^{\prime}$ to be ${\bf x}(g_{i})$ for $i\leq n$ and ${\bf x}^{\prime}(g_{i})=0$ for $i>n$, we obtain $r^{\prime}$ such that $m(r^{\prime})=0$. It can happen that $m(r)\leq\frac{1}{2}$. In this case we define $\bf x^{\prime}$} so that ${\bf x}^{\prime}(g_{i})=1$ for all $i>n$. It represents $r^{\prime}$ such that $m(r^{\prime})=1$. We conclude by the following statement.
Proposition 4.6.
The measure $m(r)$ is not a partial recursive function in any interval from $[0,1]$.
It is worth mentioning that Theorem 3.2 of [23] describes polynomial time computable real valued functions as some special limits of simple piecewise linear functions. Convergence moduli naturally appear in the definition of these limits, see Definition 3.8 of that paper. This seems slightly analogous to non-computability of $m(r)$ and non-boundedness of convergence moduli in $m_{i}(r)\to m(r)$.
5 Metric groups and amenability
Locally compact groups form the basic area of classical amenability theory. Thus analysis of computability aspects of amenability in the case of metric locally compact groups is a natural challange. On the other hand motivated by strong progress made in the last two decates in the general case of amenable topological groups (see [15] [21], [30], [31], [32]), we study the subject without the restriction of local compactness.
In this section a natural framework for computable amenability of metric groups is presented. We will see that some results from the previous sections have natural counterparts in the metric case. On the other hand, we will also describe some new issues compared to the discrete case.
5.1 Computable presentations of metric groups
We consider computably enumerable metric groups using the standard approach of computability theory. It basically corresponds to computable presentations of Polish spaces considered in [28] and [35]. See also more recent papers [26], [7], [13] where some kinds of computable presentations of continuous metric structures is considered.
We usually assume that a metric group is considered with a right-invariant metric $\leq 1$. This is a more general case compared to papers mentioned above, where bi-invariantness is usually assumed.
Definition 5.1.
Let $(G,d)$ be a metric group and $\nu:\mathbb{N}\rightarrow G$ be a function such that $\nu(\mathbb{N})$ is dense in $G$. We call the triple $(G,d,\nu)$ a numbered metric group. The function $\nu$ is called a numbering of $(G,d)$. If $g\in G$ and $\nu(n)=g$, then $n$ is called a number of $g$.
It is worth noting that a numbered metric group $(G,d,\nu)$ can be viewed together with the additional condition that $\nu(\mathbb{N})$ is a subgroup of $G$ and $\nu$ is a homomorphism from the enumerated group $(\mathbb{N},*,^{-1},1)$ introduced in Section 2.1 (i.e. a computable copy of $\mathbb{F}_{\omega}$). Indeed, the numbering $\nu$ from the definition, naturally extends to a homomorphism
$\mathbb{F}_{\omega}\to\{w(\nu(n_{1}),\ldots,\nu(n_{s}))\,|\,w(x_{1},\ldots,x_{s})\mbox{ is a group word over }x_{1},\ldots,x_{s},\ldots\}\leq G,$
which can be viewed as a numbering:
$\nu_{1}:\mathbb{N}\to\{w(\nu(n_{1}),\ldots,\nu(n_{s}))\,|\,w(x_{1},\ldots,x_{s})\mbox{ is a group word over }x_{1},\ldots,x_{s},\ldots\}$
being a homomorphism $(\mathbb{N},*,^{-1},1)\to G$. This explains why in the definition below we take this assumptions (see also the paragraph after the definition).
Definition 5.2.
  • Given an enumerated group $(\mathbb{N},\star,^{-1},1)$ we call a homomorphism $\nu:\mathbb{N}\to G$ a computably enumerable presentation of $G$ if $\nu(\mathbb{N})$ is dense in $G$ and the sets
    $\mathsf{Wrd}^{=}_{\nu}:=\{(w(n_{1},\ldots,n_{s}),w^{\prime}(\ell_{1},\ldots,\ell_{t}))\,|\,w(\bar{x})\mbox{ and }w^{\prime}(\bar{y})\mbox{ are group words and }$
    $w(\nu(n_{1}),\ldots,\nu(n_{s}))=w^{\prime}(\nu(\ell_{1}),\ldots,\nu(\ell_{t}))\mbox{ holds in }G\,\}$
    and
    $\mathsf{Wrd}^{<}_{\nu}:=\{(w(\nu(n_{1}),\ldots,\nu(n_{s})),w^{\prime}(\nu(\ell_{1}),\ldots,\nu(\ell_{t})),k):\,w(\bar{x}),w^{\prime}(\bar{y})\mbox{ are group words and }$
    $d(w(\nu(n_{1}),\ldots,\nu(n_{s})),w^{\prime}(\nu(\ell_{1}),\ldots,\nu(\ell_{t})))<\frac{1}{k}\}$
    are computably enumerable.
  • If $(G,d)$ has a computably enumerable presentation then we say that $(G,d)$ is computably enumerable.
  • If a homomorphism $\nu:(\mathbb{N},\star,^{-1},1)\to G$ is a computably enumerable presentation of $G$ and the sets $\mathsf{Wrd}^{=}$ and $\mathsf{Wrd}^{<}$ are computable, then we say that $\nu$ is a computable presentation and the group $(G,d,\nu)$ is computable.
Note that if $(G,d,\nu)$ is a numbered metric group such that the sets $\mathsf{Wrd}^{=}_{\nu}$ and $\mathsf{Wrd}^{<}_{\nu}$ are computably enumerable, then the group $(G,d,\nu_{1})$ built after Definition 5.1 is computably enumerable too. Indeed, since any word $u(\nu_{1}(m_{1}),\ldots,\nu_{1}(m_{s}))$ coincides with some $w(\nu(n_{1}),\ldots,\nu(n_{s}))$ and the latter one can be found in a computable way, any enumeration of the sets $\mathsf{Wrd}^{=}$ and $\mathsf{Wrd}^{<}$ with respect to $\nu$ effectively determines an enumeration of the corresponding set with respect to $\nu_{1}$.
Remark 5.3.
Assume that a numbered metric group $(G,d,\nu)$ is defined for an enumerated group $(\mathbb{N},\star,^{-1},1)$ with computable operations and $\nu$ is a group homomorphism. Then the group is computable if and only if the set
$\mathsf{Wrd}^{\leq}_{\nu}:=\{(w(\nu(n_{1}),\ldots,\nu(n_{s})),w^{\prime}(\nu(\ell_{1}),\ldots,\nu(\ell_{t})),k):\,k\in\mathbb{N}\cup\{\infty\}\,\,,\,w(\bar{x}),w^{\prime}(\bar{y})\mbox{ are group words and }$
$d(w(\nu(n_{1}),\ldots,\nu(n_{s})),w^{\prime}(\nu(\ell_{1}),\ldots,\nu(\ell_{t})))\leq\frac{1}{k}\}$
is computable (we identify $0=\frac{1}{\infty}$). Furthermore, it is easy to see that
the group is computable if and only if it is computably enumerable and there is an algorithm which for any $i,j\in\mathbb{N}$ and $\varepsilon\in\mathbb{Q}^{+}$ finds a rational number $\ell$ such that $d(\nu(i),\nu(j))\in[\ell,\ell+\varepsilon)$.
This shows that $d(\nu(i),\nu(j))$ is a computable real number.
Remark 5.4.
As in Section 2.1 (see Lemma 2.3) we may assume that when $(G,d,\nu)$ is computably enumerable, the operations of the enumerated group $(\mathbb{N},\star,^{-1},1)$ are computable (in fact, we just take the free group $(\mathbb{N},*,^{-1},1)$). Note that under this assumption the set
$\mathsf{MultT}:=\{(i,j,k)\,|\,\nu(i)\nu(j)=\nu(k)\}$
is computable exactly when the set $\mathsf{T}_{=}:=\{(i,j)\,|\,\nu(i)=\nu(j)\}$ is computable.
From now on we will consider computably enumerable groups such that $\nu$ is a homomorphism from the computable enumerated group $(\mathbb{N},*,^{-1},1)$ and the image of the numbering is a dense subgroup.
Remark 5.5.
It is worth noting that when a numbered metric group $(G,d,\nu)$ is computably enumerable and the set $\mathsf{T}_{=}$ is computable, the numbering $\nu$ in the statement above can be chosen to be injective.
Remark 5.6.
Assume that $G$ is a conutable discrete group. Let us consider it with respect to the $\{0,1\}$-metric $d$. The following statements are easy.
  • If $(G,\nu)$ is computably enumerable in the sense of Section 2.1, then the metric group $(G,d,\nu)$ is computably enumerable.
  • If $(G,\nu)$ is computable in the sense of Section 2.1, then $(G,d,\nu)$ is a computable presentation in the sense of Section 2.4.
5.2 Amenability and effective amenability
In order to develope computable amenability in the metric case, we apply amenability theory of topological groups developed by F.M. Schneider and A. Thom in [31]. A rough presentation of it is as follows.
Let $G$ be a topological group, $F_{1},F_{2}\subset G$ are finite and $U$ be an identity neighbourhood. Let $R_{U}$ be a binary relation defined as follows:
$R_{U}=\{(x,y)\in F_{1}\times F_{2}:yx^{-1}\in U\}.$
This relation defines a bipartite graph on $(F_{1},F_{2})$, say $\Gamma$. A matching in $\Gamma$ is an injective map $\phi:D\to F_{2}$ such that $D\subseteq F_{1}$ and $(x,\phi(x))\in R_{U}$ for all $x\in D$. A matching $\phi$ in $\Gamma$ is said to be perfect if $dom(\phi)=F_{1}$. Furthermore, the matching number of $\Gamma$ is defined to be
$\mu(\Gamma)=\mathsf{sup}(|\mathsf{dom}(\phi)||\phi\mbox{ matching in }\Gamma\}.$
By Hall’s matching theorem this value is computed as follows:
$\mu(F_{1},F_{2},U)=|F_{1}|-\mathsf{sup}\{|S|-|N_{R}(S)|:S\subseteq F_{1}\},$
where $N_{R}(S)=\{y\in F_{2}:(\exists x\in S)(x,y)\in R_{U}\}$.
The following notion will be convenient too.
Definition 5.7.
Given $\varepsilon$ and $g\in G$ we call an injective map $\phi:F\to gF$ an $\varepsilon$-matching for $gF$ if for every $g^{\prime}\in F$, $d(g^{\prime},\phi(g^{\prime}))<\varepsilon$ and the size of $\mathsf{dom}(\phi)$ is maximal for this condition.
Theorem 4.5 of [31] gives the following description of amenable topological groups.
Let $G$ be a Hausdorff topological group. The following are equivalent.
(1) $G$ is amenable.
(2) For every $\theta\in(0,1)$, every finite subset $D\subseteq G$, and every identity neighbourhood $U$, there is a finite non-empty subset $F\subseteq G$ such that
$\forall g\in D(\mu(F,gF,U)\geq\theta|F|).$
(3) There exists $\theta\in(0,1)$ such that for every finite subset $D\subseteq G$, and every identity neighbourhood $U$, there is a finite non-empty subset $F\subseteq G$ such that
$\forall g\in D(\mu(F,gF,U)\geq\theta|F|).$
It is worth noting here that when an open neighbourhood $V$ contains $U$ the number $\mu(F,gF,U)$ does not exceed $\mu(F,gF,V)$. In particular in the formulation above we may consider neighbourhoods $U$ from a fixed base of identity neighbourhoods. For example in the case of a right-invariant metric group $(G,d)$ we may take all $U$ in the form of metric balls $B_{<q}=\{x:d(1,x)<q\}$, $q\in\mathbb{Q}\cap(0,1)$. It is also clear that we can restrict all $\theta$ by rational ones. From now on we work in this case.
Let $B_{q}=\{x:d(1,x)\leq q\}$, $q\in\mathbb{Q}\cap(0,1)$. Notice that the corresponding versions of statement (2) above are equivalent for $U$ of the form $B_{<q}$ and of the form $B_{q}$. Indeed, this follows from the observation that $\mu(F,gF,B_{<q})\leq\mu(F,gF,B_{q})$ and $\mu(F,gF,B_{q})\leq\mu(F,gF,B_{<r})$ for $q<r$.
The following observation is an important point used in our approach.
$\bullet$ It is well-known that a topological group is amenable if and only if it contains a dense amenable subgroup. Furthermore, then every dense subgroup is amenable. In particular, the formulation of the Schneider-Thom theorem given above still holds if in conditions (2) and (3) the group $G$ is replaced by a fixed dense subgroup.
Remark 5.8.
When $G$ is a contable group considered with respect to the $\{0,1\}$-metric, the set $U$ appearing in (2) and (3) can be taken $U=\{e\}$. Note that in this case the number $\mu(F_{1},F_{2},U)$ is just $|F_{1}\cap F_{2}|$. In particular, taking $\theta=\frac{n-1}{n}$ the condition that the subset $F\subset G$ satisfies
$\forall g\in D(\mu(F,gF,U)\geq\theta|F|)$
just means that $F$ is $\frac{1}{n}$-Følner with respect to $D$.
This remark suggests calling $F$ with
$\forall g\in D(\mu(F,gF,B_{<q})\geq\frac{n-1}{n}|F|)$
to be $\frac{1}{n}$-Følner with respect to $D$ and $q$. Now we will say that a sequence $(F_{j})_{j\in\mathbb{N}}$ of non-empty finite subsets of $G$ is a Følner sequence if for every $g\in G$ the following condition holds:
$\lim\limits_{n\rightarrow\infty}\mu(F_{n},gF_{n},B_{<\frac{1}{n}})=1.$ (8)
It is easy to see that existence of Følner sets for all $n$ and $D$ is equivalent to existence of a Følner sequence. In fact, this is the theorem of Schneider and Thom stated above. It can be viewed as the metric version of Følner condition of amenability.
We can now formalize computable amenability of numbered metric groups. Assume that $(G,d,\nu)$ is a numbered, right-invariant metric group such that $\nu(\mathbb{N})$ is a dense subgroup of $G$. The situation that $G=\nu(\mathbb{N})$ and $d$ is the $\{0,1\}$-metric, is possible.
When $D\subset_{fin}\mathbb{N}$ and $m,n\in\mathbb{N}$, let us denote by $\mathfrak{F}\o l_{G,\nu(D),m}(n)$ the family of all finite subsets $F\subset\mathbb{N}$ which satisfy
$\forall e\in D\,\,(\mu(\nu(F),\nu(e)\nu(F),B_{<\frac{1}{m}}(1))\geq\frac{n-1}{n}|\nu(F)|).$
In the case of the Schneider-Thom version of Følner’s condition of amenability, we again consider two types of effectiveness: they correspond to ones from Section 3.
Definition 5.9.
The numbered metric group $(G,d,\nu)$ is $\Sigma$-amenable, if there is an algorithm which for all triples $(m,n,D)$ where $m,n\in\mathbb{N}$ and $D\subset_{fin}\mathbb{N}$, finds a set $F\subset_{fin}\mathbb{N}$ having $F^{\prime}\subseteq F$ with $F^{\prime}\in\mathfrak{F}\o l_{G,\nu(D),m}(n)$.
Remark 5.10.
In the case when the group $(G,d,\nu)$ is discrete with the $\{0,1\}$-metric we arrive at the formulation that it is $\Sigma$-amenable if there exists an algorithm which for all pairs $(n,D)$, where $n\in\mathbb{N}$ and $D\subset_{fin}\mathbb{N}$, finds a set $F\subset_{fin}\mathbb{N}$ with $\nu(F^{\prime})\in\mathfrak{F}\o l_{G,\nu(D)}(n)$ where $F^{\prime}\subseteq F$. This is exactly Definition 3.1.
Definition 5.11.
The numbered metric group $(G,d,\nu)$ is computably amenable, if there is an algorithm which for all quadrangles $(\ell,m,n,D)$ where $\ell,m,n\in\mathbb{N}$ and $D\subset_{fin}\mathbb{N}$, finds a set $F\in\mathfrak{F}\o l_{G,\nu(D),m}(n)$ with $|F|=|\nu(F)|$, together with an assignament $(i,j)\to q$ where $i,j\in F$, $q\in\mathbb{Q}^{+}$ and $d(\nu(i),\nu(j))\in[q,q+\frac{1}{\ell})$.
Remark 5.12.
It is easy to see that in the discrete case the group $(G,d,\nu)$ is computably amenable if and only if it satisfies Definition 3.2.
The discussion after the formulation of the Schneider-Thom theorem implies that for a numbered metric group, $\Sigma$-amenability implies amenability of $G$. It is clear that computable amenability imples $\Sigma$-amenability. Furthermore, $\Sigma$-amenability implies that the Følner function is subrecursive.
5.3 Effective amenability of computable metric groups
The following observation shows that amenable groups which have good computable presentation have computable Følner sets.
Proposition 5.13.
Assume that a computably enumerable metric group $(G,d,\nu)$ has decidable equality relation: the set $\mathsf{T}_{=}=\{(i,j)\,:\nu(i)=\nu(j)\}$ is computable. Then amenability of $G$ implies that $(G,d,\nu)$ has computable Følner sets, which means the following property:
there is an algorithm which for all triples $(m,n,D)$ where $m,n\in\mathbb{N}$ and $D\subset_{fin}\mathbb{N}$, finds a set $F\in\mathfrak{F}\o l_{G,\nu(D),m}(n)$.
Before the proof we give several remarks.
Remark 5.14.
(1) The assumptions of the first sentence of this proposition imply that the set $\mathsf{Wrd}^{=}$ is computable. To see this use the assumption that the multiplication * on $\mathsf{N}$ is a computable function.
(2) It is clear that when $(G,d,\nu)$ has computable Følner sets, it is $\Sigma$-amenable. On the other hand computable amenability implies having computable Følner sets.
(3) Note that in the discrete case $(G,d,\nu)$ has computable Følner sets if there exists an algorithm which, for all pairs $(n,D)$, where $n\in\mathbb{N}$ and $D\subset_{fin}\mathbb{N}$, finds a finite set $F\subset\mathbb{N}$ such that $\nu(F)\in\mathfrak{F}$$l_{G,\nu(D)}(n)$.
Proof.
Let us fix an enumeration of all quadrangles of the form $(m,n,D,F)$ where $m,n\in\mathbb{N}$ and $D,F\subset_{fin}\mathbb{N}$. The following procedure, denoted below by $\hat{\vartheta}(m,n,D,F)$, determines quadrangles satisfying the condition of the proposition.
For an input $(m,n,D,F)$ let $F_{0}\subseteq F\cup D*F$ be a set representing the $\mathsf{T}_{=}$-classes in $F\cup E*F$. Let us fix an enumeration of the set
$\mathsf{T}^{<}_{F_{0}}=\{(n_{1},n_{2},q):\,n_{1},n_{2}\in F_{0}\,,\,q\in\mathbb{Q}\cap(0,1)\,,\,d(\nu(n_{1}),\nu(n_{2}))<q\}.$
After the $m$-th step of this enumeration we obtain a set of restrictions on $d$ in $F_{0}$. Having this we verify whether these restrictions enforce the condition
$\forall g\in D\,\,(\mu(\nu(F_{0}\cap F),\nu(g)\nu(F_{0}\cap F),B_{<\frac{1}{m}}(1))\geq\frac{n-1}{n}|F_{0}\cap F|)$
If this is not the case we make the next step. Note that the number $\mu(\nu(F_{0}\cap F),\nu(g)\nu(F_{0}\cap F),B_{<\frac{1}{m}}(1))$ does nor decrease. Indeed, since the distances are becoming smaller, new matchings can occur, but the matchings which were already found, can only increase to larger ones. We stop when the above inequality holds. Note that if in the numbered metric group $(G,d,\nu)$ the condition
$\forall g\in D\,\,(\mu(\nu(F),\nu(g)\nu(F),B_{<\frac{1}{m}}(1))\geq\frac{n-1}{n}|\nu(F)|)$
holds, then it would be recognized at some step of the procedure $\hat{\vartheta}(m,n,E,F)$. Under the notation above, then the set $F_{0}\cap F$ would serve as the corresponding Følner set.
The algorithm for Følner sets of $(G,d,\nu)$ looks as follows. Having an input $(m,n,D)$ we enumerate all finite $F$ and for each of them start the procedure $\hat{\vartheta}(m,n,E,F)$. By amenability of $(G,d)$ for some $F$ such a procedure would give the result. ∎
The authors think that the statement of Proposition 5.13 can not be strenthenned to the condition of computable amenability. However, at the moment we do not have any counterexample. The following theorem is a metric version of Theorem 3.6. This is Theorem 3 from Introduction.
Theorem 5.15.
Let $(G,d,\nu)$ be a computably enumerable metric group. The following conditions are equivalent:
  1. (i)
    $(G,d,\nu)$ is amenable and computable;
  2. (ii)
    $(G,d,\nu)$ is computably amenable (Definition 5.11).
Proof.
[…]
Fix $n_{1},n_{2}$. Let $D$ be the set $\{n_{1},n_{2}\}$. We apply the algorithm for (ii) to some $(\ell,m,n,D)$ where $\frac{2}{\ell}\leq\varepsilon$ , $m>\frac{4}{\varepsilon}$ and $\frac{2}{3}\leq\frac{n-1}{n}$. Let $F\subset\mathbb{N}$ be a set which is the output of the algorithm and let $\Sigma^{\ell}_{F}=\{(f,f^{\prime},q)\,|\,f,f^{\prime}\in F,\,q\in\mathbb{Q}^{+},d(\nu(f),\nu(f^{\prime}))\in[q,q+\frac{1}{\ell})\}$.
For each $i\in\{1,2\}$ we define $\Sigma_{i}\subseteq\{(f,f^{\prime})\,|\,d(\nu(n_{i})\nu(f),\nu(f^{\prime}))\leq\frac{1}{m},f,f^{\prime}\in F\}$ by the following procedure. Having $f,f^{\prime}\in F$ apply the algorithm of enumeration of the set $\mathsf{Wrd}^{\leq}_{\nu}$ for verification if
$d(\nu(n_{i}*f),\nu(f^{\prime}))<\frac{1}{m}\,,\,i=1,2.$
When we get a confirmation of this inequality, we extend the corresponding $\Sigma_{i}$ by $(f,f^{\prime})$. We apply it simultaneously to each pair $(f,f^{\prime})$. Since
$\forall g\in D\,\,(\mu(\nu(F),\nu(g)\nu(F),B_{<\frac{1}{m}}(1))\geq\frac{n-1}{n}|\nu(F)|)$
and $|F|=|\nu(F)|$, there is a step of these computations when $\Sigma_{1}\cup\Sigma_{2}$ confirms the existence of matchings witnessing this inequality. Having this, we stop the procedure.
By the choice of $n$ there are pairs $(f,f^{\prime})\in\Sigma_{1}$ and $(f,f^{\prime\prime})\in\Sigma_{2}$. Let $(f^{\prime},f^{\prime\prime},q)\in\Sigma^{\ell}_{F}$. Then we have that $d(\nu(n_{1})\nu(f),\nu(n_{2})\nu(f))\in[q-\frac{1}{m},q+\frac{1}{\ell}+\frac{1}{m})$. Since $d$ is right invariant we see
$d(\nu(n_{1}),\nu(n_{2}))\in[q-\frac{1}{m},q+\frac{1}{\ell}+\frac{1}{m}).$
In particular $q-\frac{1}{m}$ serves as the requred $q_{0}$. ∎
As in the case of Theorem 3.6 we have the following interesting observation.
Corollary 5.16.
Let $(G,d,\nu)$ be a computably enumerable, amenable group. If for some $n\geq 3$ there exists an algorithm, which for every $\varepsilon>0$ and $D\subset_{fin}\mathbb{N}$ with $|D|=2$, finds a set $F\subset_{fin}\mathbb{N}$ such that $F\in F\o l_{G,\nu(D),\frac{5}{\varepsilon}}(n)$ and $|F|=|\nu(F)|$, together with an assignment as in the formulation of Definition 5.11 for $\ell\geq\frac{2}{\varepsilon}$, then $(G,d,\nu)$ is computable.
The following corollary corresponds to Theorem 3.8.
Corollary 5.17.
Let $(G,d,\nu)$ be a computable group. Then the following conditions are equivalent:
  1. (i)
    $(G,d)$ is amenable;
  2. (ii)
    $(G,d,\nu)$ is computably amenable;
  3. (iii)
    $(G,d,\nu)$ has computable Følner sets;
  4. (iv)
    $(G,d,\nu)$ is $\Sigma$-amenable.
Proof.
[…]
Comparing this corollary with Theorem 3.8 the reader observes that we do not include here any statement concerning Reiter’s functions. At the moment the authors do not have any metric version of Section 3.1. This looks as a non-trivial task.
Let us consider Følner sequences in computable metric groups . Let $(G,d,\nu)$ be a computable metric group. By Remarks 5.5 and 5.3 we may assume that the function $\nu$ is injective. Therefore we identify the set $\nu(\mathbb{N})$ with $\mathbb{N}$ and subsets $F$ of $\mathbb{N}$ with $\nu(F)\subset G$.
As in Section 4 an effective Følner sequence of the group $(G,d,\nu)$ is an effective sequence $(n_{j})_{j\in\mathbb{N}}$ where each $n_{j}$ is a Gödel number of some $F_{j}$, such that $(F_{j})_{j\in\mathbb{N}}$ is a Følner sequence.
By Theorem 5.15, amenability of $(G,d,\nu)$ is equivalent to computable amenability. This is also equivalent to existence of an effective Følner sequence. Indeed, apply the argument given in beginning of Section 4.1.
Let $\varphi(x,y)$ be a universal recursive function, and $\varphi_{x}(y)=\varphi(x,y)$ be the recursive function with the number $x$. We identify effective Følner sequences with numbers of recursive functions which produce these sequences. The set of these numbers will be denoted by $\mathfrak{F}_{seq}(G,d,\nu)$. The description of this set in the arithmetical hierarchy (see Theorem 4.1) has the following counterpart.
Let $(G,d,\nu)$ be a computable group. The set of all effective Følner sequences of $(G,d,\nu)$ belongs to the class $\Pi^{0}_{3}$.
Indeed, it is straightforward that $m\in\mathfrak{F}_{seq}(G)$ if and only if the following formula holds:
$(\phi(m,y)$ $\text{ is a total function})\wedge(\forall g\in\nu(\mathbb{N}))(\forall n)(\exists l)(\forall k)\Big{(}k>l\wedge(\phi(m,k)=f)$
[…]
Given number $f$ the inequality $\mu(F,gF,B_{<\frac{1}{n}}(1))\geq\frac{n-1}{n}|F|$ can be verified effectively. Since the set of numbers of all total functions belongs to the class $\Sigma_{2}^{0}$ it is easy to see that the set of all $m$ which satisfy (9) is a $\Pi_{3}^{0}$ set.